State the assumptions that are made in the kinetic theory of gases and derive an expression for the pressure exerted by a gas which conforms to these assumptions, in terms of density and the root mean square velocity of its molecles
1
Expert's answer
2021-02-10T13:13:22-0500

Answer

According to KTG assumption are as below

  • The space-volume to molecules ratio is negligible as physics.
  • There is no force of attraction between the molecules at the normal temperature and pressure.
  • The force of attraction between the molecules builds when the temperature decreases and the pressure increases.
  • There is a large space between the molecules resulting in continuous motion KTG.
  • The free movement of molecules results in a collision which is perfectly elastic.
  • The molecules have kinetic energy due to random movement. But the average kinetic energy of these molecules differs with temperaturE in MOTION.
  • Molecules exert pressure on the walls of the container.
  • derivation of expression-

As pressure

"P=\frac{m_T N}{3VN}\Sigma v^2\\=\frac{M_T}{3VN}\Sigma v^2\\ =\frac{\rho v_{rms}^2}{3}"